OpenStudy (haleyelizabeth2017):

Honors Algebra 2A~Question and options in comments :)

OpenStudy (haleyelizabeth2017):

What is the solution to the following system? -2x-y=1 -4x-2y=-1 I think no solutions...

OpenStudy (anonymous):

Get them in slope intercept form first

OpenStudy (vlery):

-.5 and .25 all you have to do is graph it and see where it crosses the x axis

OpenStudy (haleyelizabeth2017):

the options are (3, 8), (-3, -8) no solutions or infinant solutions. I did substitution to find x and y and the x's canceled when I was solving for x and so no solutions?

OpenStudy (haleyelizabeth2017):

@SydtheKid913

OpenStudy (anonymous):

When you put it into slope-intercept you get: \[y=-2x-1\] and \[y=-2x+\frac{ 1 }{ 2 }\] They have the same exact slope. This means they are parallel and therefore have no solution

OpenStudy (haleyelizabeth2017):

Thank you so much!

OpenStudy (anonymous):

You're welcome

OpenStudy (haleyelizabeth2017):

\(\color{#F52887}{\Huge\ddot\smile}\)